OpenStudy (anonymous):

write an equation for the line perpendicular to y=-3x+6 that contains(7,-4)

OpenStudy (anonymous):

y +4 = 1/3(x-7) y = 1/3x +5/7

OpenStudy (anonymous):

y = 1/3x -19/7